# Minimum Cost Loop Nests for Contraction of a Sparse Tensor with a Tensor Network

### v1

@@ Abstract.
Sparse tensor decomposition and completion are common in numerous applications, ranging from machine learning to computational quantum chemistry. Typically, the main bottleneck in optimization of these models are contractions of a single large sparse tensor with a network of several dense matrices or tensors (SpTTN). Prior works on high-performance tensor decomposition and completion have focused on performance and scalability optimizations for specific SpTTN kernels. We present algorithms and a runtime system for identifying and executing the most efficient loop nest for any SpTTN kernel. We consider both enumeration of such loop nests for autotuning and efficient algorithms for finding the lowest cost loop-nest for simpler metrics, such as buffer size or cache miss models. Our runtime system identifies the best choice of loop nest without user guidance, and also provides a distributed-memory parallelization of SpTTN kernels. We evaluate our framework using both real-world and synthetic tensors. Our results demonstrate that our approach outperforms available generalized state-of-the-art libraries and matches the performance of specialized codes.
@@
conference: ; ; copyright: noneccs: Software and its engineering General programming languagesccs: Social and professional topics History of programming languages
@@ 1. Introduction
Prior works with focus high-performance tensor decomposition and completion have introduced efficient and parallel implementations for many SpTTN kernels (Smith et al., 2015; Kaya and Uçar, 2015; Li et al., 2017; Phipps and Kolda, 2019; Choi et al., 2018; Li et al., 2018). Most of these works focus specifically on one or two kernels needed for a particular tensor decomposition, e.g., the matricized Khatri-Rao product (MTTKRP) for CP decomposition (Choi and Vishwanathan, 2014; Kang et al., 2012; Blanco et al., 2018) or the tensor times matrix chain (TTMc) kernel for Tucker (Smith and Karypis, 2017; Oh et al., 2018)). Even for a single decomposition, different algorithms often rely on different SpTTN kernels (Singh et al., 2022). By developing algorithms and libraries for arbitrary SpTTN kernels, we provide functionality for contraction arising (e.g., as a result of a gradient calculation, described in Section 3) from any decomposition/network consisting of dense tensors.
The main challenge in implementation of an SpTTN kernel is finding the most efficient loop nest. In line with prior work (Kjolstad et al., 2017), we represent such loop nest as a trees, in which each vertex is a loop and its descendants are the loops contained within it. In Section 4, we show how to enumerate all loop nests (assuming fusion is done wherever possible) for a given SpTTN. Since each loop order for any pair of contracted tensors yields a distinct loop nest, this size of this space grows factorially in the loop nest depth $m$ and exponentially in the number of tensors $N$. We provide a dynamic programming algorithm to find a cost-optimal loop nest with substantially lower cost, namely $O(N^{3}2^{m}m)$ instead of $O((m!)^{N})$. We state the algorithm for a general cost function that can be decomposed according to the loop nest tree structure, then provide specific cost functions to minimize buffer size and cache complexity.
@@ 2. Background
(b) Vertices $i$ and $j$ are fused across the two pairwise contractions. The order of the intermediate tensor $\mathcal{X}$ is reduced by 2 since $i$ and $j$ are no longer buffered.
(c) Vertices $i$, $j$ and $s$ are fused across the two pairwise contractions. The loop_order is (i,j,s,k,r). The intermediate tensor $X$ is a scalar as none of the indices are buffered.
(d) Contraction sequence where $\mathcal{U}$ and $\mathcal{V}$ are contracted first, followed by $\mathcal{T}$. None of the vertices can be fused.
Figure 1. Iteration graphs illustrating various loop nests for computing an order 3 TTMc kernel. Sparse loops are shown as dotted vertices.
2.1. Tensor Notation
We use calligraphic letters to denote tensors, e.g., $\mathcal{T}$. An order $N$ tensor corresponds to an $N$-dimensional array. We denote elements of tensors in parenthesis, e.g., $\mathcal{T}(i,j,k,l)$ for an order 4 tensor $\mathcal{T}$. We do not use distinct notation for vectors and matrices, but instead treat them as order 1 and order 2 tensors, respectively. A tensor contraction specifies a way to multiply two or more tensors, by assigning an index to each dimension of the tensors. The indices that do not appear in the output tensor are considered to be summed (contracted). For example, matrix-matrix multiply that is contracted over index $k$ is denoted as $\mathcal{A}(i,j)=\sum_{k}\mathcal{B}(i,k)\cdot\mathcal{C}(k,j)$. We use capitalized letters to denote the dimensions of the respective indices. For example, the dimension of index $i$ in $\mathcal{A}(i,j)$ is denoted as $I$.
2.2. Tensor Sparsity and Sparse Storage
One of the most common ways to store sparse tensors is the Compressed Sparse Fiber (CSF) format (Smith and Karypis, 2015), which is a generalization of Compressed Sparse Row (CSR)/Compressed Sparse Column (CSC) format for matrices. CSF can be viewed as a tree with height equal to the order of the tensor. Each index of the tensor is associated with a level in the tree. Every level except the last stores a list of nonzero index values of that level and a list of pointers to the nonzero index values of the next level. The last level (the set of leaves) stores a list of nonzero index values and the tensor data represented by that branch of the tree. Traversing the CSF tree representation of a tensor amounts to iterating over the tensor dimensions in the order that they are stored (from root to leaves).
We refer to the total number of nonzero elements of a tensor $\mathcal{T}$ as $\operatorname{nnz}(\mathcal{T})$. For a sparse tensor $\mathcal{T}$ with $d$ dimensions of size $I_{1},\ldots,I_{d}$, we represent the number of non-zeroes in the $k$th level of the CSF tree for $\mathcal{T}$ (with the first index being at the root) as $\operatorname{nnz}^{(I_{1}\cdots I_{k})}(\mathcal{T})$. Equivalently, this nonzero count may be obtained by considering the number of nonzeros in a reduced tensor obtained by summing away the remaining modes, i.e., $\operatorname{nnz}^{(I_{1}\cdots I_{k})}(\mathcal{T})=\operatorname{nnz}(\mathcal{S})$, where
$\mathcal{S}(i_{1},\ldots,i_{k})=\sum_{i_{k+1},\ldots,i_{d}}|\mathcal{T}(i_{1},\ldots,i_{d})|.$
2.3. Tensor Decomposition and Completion Algorithms
Tensor decomposition (Kolda and Bader, 2009) and completion (Smith et al., 2016) refer to the problem of decomposing a tensor into a combination of smaller tensors and estimating missing or incomplete values in a tensor, respectively. These techniques are often used for feature extraction, compression, or when the tensor data is partially observed or noisy. Tensor decomposition and completion can be seen as generalization of matrix decomposition and completion, respectively.
2.4.1. Unfactorized Contraction
To compute a kernel, we can iterate over all the indices and simultaneously contract all the input tensors in the innermost loop. We refer to this approach as unfactorized. This unfactorized loop nest has a depth equal to the number of unique indices. For example, consider an order 3 TTMc kernel in Equation 2, where each of the factor matrices have one unshared index with the tensor, and have no shared indices among them. The unfactorized approach for the kernel computation has to iterate over all the 3 tensor indices and each of the unique factor matrix indices resulting in a depth of five. The number of operations is $3\operatorname{nnz}(\mathcal{T})\cdot R\cdot S$ to leading order. In compiler driven frameworks such as TACO (Kjolstad et al., 2017) and COMET (Tian et al., 2021), the schedule generated by default is unfactorized.
The unfactorized approach is optimal in cost for computing certain kernels. For example, the MTTKRP kernel in Equation 1 can be computed using the unfactorized approach with an optimal loop depth of four. But this approach is asymptotically suboptimal for many other kernels, such as the TTMc kernel.
@@ 10        $\mathcal{S}$[i,r,s] += $\mathcal{X}$[s] * $\mathcal{U}$[j,r]
2.4.2. Pairwise Contraction
A kernel can be computed with minimal asymptotic complexity (loop depth) by contracting the input tensors pairwise. This approach is typically used in libraries designed for dense tensor contractions, such as CTF (Solomonik et al., 2014), Tensor Computation Engine (TCE) (Baumgartner et al., 2005), and DEinsum (Ziogas et al., 2022). We first pick any two input tensors to contract, then contract its output with an input tensor that is not yet contracted. We repeat this process until all the tensors are contracted. We refer to this approach as pairwise contraction. The loop depth for TTMc can be reduced by contracting the tensors pairwise. For example, one way in which the tensors can be contracted pairwise is to first contract $\mathcal{T}$ with $\mathcal{V}$, and then its result with $\mathcal{U}$. Equation 5 shows the 2 pairwise contractions $((\mathcal{T}\cdot\mathcal{V})\cdot\mathcal{U})$,
$\mathcal{X}(i,j,s)=\sum_{k}\mathcal{T}(i,j,k)\cdot\mathcal{V}(k,s),$
(5) $\mathcal{S}(i,r,s)=\sum_{j}\mathcal{X}(i,j,s)\cdot\mathcal{U}(j,r),$
where $\mathcal{X}$ is an intermediate tensor. Each such pairwise contraction has an independent loop nest as shown in Listing LABEL:lst:ttmc_pairwise. For this TTMc pairwise contractions, both the loop nests have a depth of 4, and the computational cost is $2\operatorname{nnz}(\mathcal{T})\cdot S+2\operatorname{nnz}^{(IJ)}\cdot S\cdot R$ to leading order.
Even though an unfactorized approach for computing the MTTKRP kernel (Equation 1) has an optimal loop depth, up to a third of the operations can be eliminated by using pairwise contraction. The unfactorized approach requires $3\cdot\operatorname{nnz}(\mathcal{T})\cdot A$ scalar operations, while the pairwise approach requires $2\operatorname{nnz}^{(IJK)}(\mathcal{T})\cdot A+2\operatorname{nnz}^{(IJ)}\cdot A$ operations, where $A$ is the size of the dimension $a$. This cost is achieved by contracting the tensors $\mathcal{T}$ and $\mathcal{C}$ pairwise, and then contracting the result with $\mathcal{B}$,
$\mathcal{X}(i,j,a)=\sum_{k}\mathcal{T}(i,j,k)\cdot\mathcal{C}(k,a),$
(6) $\mathcal{A}(i,a)=\sum_{j}\mathcal{X}(i,j,a)\cdot\mathcal{B}(j,a),$
where $\mathcal{X}$ is an intermediate tensor.
For contractions involving only dense tensors, the pairwise approach can provide an optimal schedule. But for sparse tensors, whose dimensions are often large, this approach can lead to unmanageable memory requirements for storing dense intermediate tensors. In both TTMc and MTTKRP kernels, the intermediate tensor $\mathcal{X}$ is prohibitively large if stored in dense format (is of size $O(I\cdot J\cdot S)$ and $O(I\cdot J\cdot A)$, respectively). While $\mathcal{X}$ is sparse (with $\operatorname{nnz}(\mathcal{X})=\operatorname{nnz}^{(IJ)}(\mathcal{T})R$ for MTTKRP and similar for TTMc), it is not efficiently represented in CSF, since one of the dimensions is dense. In practice, pairwise contraction with sparse storage of such an intermediates has been observed to be much slower than hand-tuned or even unfactorized implementations for SpTTN kernels (Singh et al., 2022).
2.4.3. Factorize-and-Fuse
The size of the intermediate tensors can be reduced by loop fusion. Loop nests that share a common index can be nested together with an outer loop that iterates over the shared index. The loop nests that compute the pairwise contractions in Equation 5 can be fused together as shown in Listing LABEL:lst:ttmc_fusion. We refer to this approach as factorize-and-fuse. The contractions are nested in a single loop over $i$, and hence the index is not buffered. Similarly, index $j$ is fused across the contractions reducing the order of $\mathcal{X}$ by one more. The computation cost remains the same as in the pairwise case (in fact, the same set of operations is computed). The dimension of the intermediate tensor $\mathcal{X}$ is reduced from $I\times J\times S$ to $S$. Hand-tuned implementations of SpTTN kernels incorporate loop fusion in some form.
Specialized libraries for some of these kernels use a similar approach in their hand-tuned implementations (Smith et al., 2015; Kaya and Uçar, 2015; Li et al., 2017; Phipps and Kolda, 2019; Choi et al., 2018; Li et al., 2018; Kang et al., 2012; Smith and Karypis, 2017). For example, SPLATT (Smith et al., 2015) achieves reduced arithmetic complexity by employing pairwise contractions (Equation 6) and reduced intermediate tensor size by fusing the loop nests that compute these pairwise contractions.
3. SpTTN Kernels
The optimization methods generally leverage all or parts of the gradient of the residual error norm, which yields a contraction of the sparse tensor, with subsets (all but one of the) tensors from the decomposition i.e. the gradient of $f$ will be
$\nabla f(\mathcal{A}_{1},\dots,\mathcal{A}_{n})=[-2\mathcal{T}\cdot\tilde{\mathcal{T}}(\varnothing,\mathcal{A}_{2},\dots,\mathcal{A}_{n})+$
$\tilde{\mathcal{T}}(\varnothing,\mathcal{A}_{2},\dots,\mathcal{A}_{n})\cdot\tilde{\mathcal{T}}(\varnothing,\mathcal{A}_{2},\dots,\mathcal{A}_{n});$
$-2\mathcal{T}\cdot\tilde{\mathcal{T}}(\mathcal{A}_{1},\varnothing,\mathcal{A}_{3},\dots,\mathcal{A}_{n})+$
$\tilde{\mathcal{T}}(\mathcal{A}_{1},\varnothing,\mathcal{A}_{3},\dots,\mathcal{A}_{n})\cdot\tilde{\mathcal{T}}(\mathcal{A}_{1},\varnothing,\mathcal{A}_{3},\dots,\mathcal{A}_{n});\cdots].$
The terms involving $\mathcal{T}$ are cost-dominant. Similarly, when computing the residual error ($\rho$) for tensor completion, which is often employed, e.g., in coordinate descent methods, the terms involving the sparse tensor are cost-dominant.
$\rho=T-\Omega\ast\tilde{T}(A_{1},\ldots,A_{n}),$
where $\ast$ is the Hadamard (pointwise product), the sparse tensor $\Omega$ represents the set of observed entries in the input tensor $\mathcal{T}$ and $\tilde{\mathcal{S}}$ is the output tensor obtained by contracting $\Omega$ with a network of dense factors.
In general, we define an SpTTN kernel as a contraction of a sparse tensor with a set of dense tensors resulting in an output with a dense representation or a sparse tensor with the same sparsity as the sparse input tensor. Hence, in any SpTTN a subset of the indices in the contraction has a fixed/known sparsity pattern (associated with the input sparse tensor), while the remaining indices iterate only over dense tensors. We generally assume the dense tensors involved in the SpTTN are fairly small (in comparison to the input sparse tensor).
3.1. Loop Nests and Loop Nest Forests
A pairwise contraction approach involves contracting two tensors at a time. We refer to these two tensor contractions as contraction terms. We refer to the order in which the terms are executed as the contraction path. It is a sequence of contraction terms and is valid if we can obtain the output tensor by executing the contraction terms in the order specified by it.
The computation of a tensor contraction generally involves loop nests of some form. Any loop nest can be represented by an ordered tree or forest, each vertex of which is a loop, and its ordered children are the loop nests contained directly in that loop. Each leaf corresponds to a contraction term, e.g., $\mathcal{T}_{ijk}\cdot\mathcal{V}_{ks}\!\rightarrow\!\mathcal{X}_{ijs}$. A similar representation is used in TACO (Kjolstad et al., 2017). Loop nests can also be represented as expression trees as used in (Bibireata et al., 2004) for dense tensor contractions.
In a valid loop nest forest, all indices in a contraction term should be loop indices on the path between the corresponding leaf and the tree root, and the path should contain no additional or repeated indices. We refer to this order of loop indices as the index order of the contraction term. If a vertex has multiple leaves in its subtree, the loop associated with that vertex contains all the contraction terms in that subtree. The leaves of the loop nest tree should define a valid contraction path for computing a kernel. We say a loop nest tree is fully-fused if no vertex has two consecutive children that correspond to the same index. A loop nest forest is fully-fused if adding a dummy vertex and connecting it to all roots in the forest yields a fully-fused loop nest tree. We refer to a tree with a single leaf as a path graph.
Figure 1 shows loop nest forests for TTMc with an order 3 sparse tensor (Equation 2). A pairwise contraction of the kernel has two path graphs (independent trees, one for each contraction term), and the contraction path gives the order in which these terms are computed. This loop nest forest is equivalent to the loop nest presented in Listing LABEL:lst:ttmc_pairwise. Similarly, loop nest for an unfactorized contraction of the kernel can be represented by a path graph. The leaf corresponds to contracting the input tensors unfactorized to obtain the output tensor. The factorize-and-fuse approach in Listing LABEL:lst:ttmc_fusion yields a fully-fused tree as shown in Fig 0(b).
3.2. Intermediate Tensors
3.3. Contraction Path and Index Order
The contraction path affects the asymptotic complexity (loop depth) and memory requirements (intermediate tensor sizes) of the computation. For example, consider the various ways to compute the TTMc kernel as shown in Figure 1. In one of the chosen contractions paths, tensors $\mathcal{T}$ and $\mathcal{V}$ are contracted first and the result is then contracted with $\mathcal{U}$. The computation has a maximum loop depth of 4. The loop nest forest in Figure 0(a) can be fused to reduce the intermediate tensor size as show in Figure 0(b). In Figure 0(d), we pick a different contraction path where tensors $\mathcal{U}$ and $\mathcal{V}$ are contracted first and then the result is contracted with $\mathcal{T}$. This path results in a maximum loop depth of 5. Also, since none of the vertices can be fused, the intermediate tensor $\mathcal{X}$ is of order 4. The corresponding loop nest is shown in Listing LABEL:lst:ttmc_contraction_order
Similarly, for a given contraction path, the ordering of vertices in the path graphs before fusing them, affects the intermediate tensor sizes and other cost metrics of interest. In the previous example of the TTMc kernel, consider the iteration graph in Figure 0(a) and its fully-fused variant in Figure 0(b). Indices $i$, $j$ and $s$ are common across the two trees in the iteration graph. We are able to fuse vertices $i$ and $j$ but not $s$. To fuse vertex $s$, the vertices $k$ and $s$ in the path graph corresponding to the first term should be interchanged. This fusion, in turn, requires the index order to be $(i,j,s,k)$ in the first path graph. A fully-fused iteration graph with such an index order is shown in Figure 0(c) and its corresponding loop nest in Listing LABEL:lst:ttmc_loop_order. Since a single loop of $s$ iterates over both the terms, the term which generates the intermediate tensor and the term which uses it, $s$ is not buffered. The intermediate tensor is now only a scalar value. Note that since the ordering of indices do not affect the height of an iteration graph, the asymptotic complexity for a contraction path remains the same for any choice of index order.
1T_csf = CSF(${\mathcal{T}\_ijk}$)
2// $j$ and $k$ in the first loop nest are treated as dense loops
[…]
4for j in range(J):
5  for k in range(K):
6    for r in range(R):
7      for s in range(S):
8        $\mathcal{X}$[j,k,r,s] += $\mathcal{U}$[j,r] * $\mathcal{V}$[k,s]
9for (i,T_ijk) in T_csf:
10  for (j,T_ij) in T_i:
11    for (k,T_ijk) in T_ij:
12      for r in range(R):
13        for s in range(S):
14          $\mathcal{S}$[i,r,s] += $\mathcal{X}$[j,k,r,s] * t_ijk
Listing 3: TTMc kernel computed using the factorize-and-fuse approach. Tensors $\mathcal{U}$, $\mathcal{V}$ are contracted first to produce an intermediate tensor $\mathcal{X}$. The main tensor $\mathcal{T}$ is then contracted with $\mathcal{X}$ to compute the output tensor $\mathcal{S}$.
@@ 9        $\mathcal{S}$[i,r,s] += $\mathcal{X}$ $\mathcal{U}$[j,r]
Listing 4: TTMc kernel computed using the factorize-and-fuse approach. A single loop nest of $i$,$j$,$s$ is used to iterate over both the pairwise contractions. Index orders in terms $\mathcal{T}\cdot\mathcal{V}\rightarrow\mathcal{X}$ and $\mathcal{X}\cdot\mathcal{U}\rightarrow\mathcal{S}$ are $(i,j,s,k)$ and $(i,j,s,r)$, respectively.
4. Algorithm to Find Optimal SpTTN Kernels
To determine an efficient loop nest for an SpTTN kernel, we seek an approach to enumerate fully-fused trees and efficient algorithms to find an optimal tree for simple cost metrics. Enumeration enables autotuning, but for analytic metrics of performance such as buffer size, more efficient search schemes are possible. Different contraction paths yield different fully-fused loop nests, hence we focus our attention to enumeration and search of loop nests for a particular contraction path. Our efficient search algorithm employs dynamic programming, after decoupling order of terms from the tree structure.
4.1. Enumeration of Loop Nests
We seek to find cost-optimal loop nests for a given SpTTN kernel by enumerating only fully-fused loop nest forests and restrict our attention to dense multidimensional buffers (intermediate tensors). We decouple the enumeration into two steps: (1) enumeration of valid contraction paths for a given set of tensors and (2) enumeration of index orders in the path graphs for a given contraction path.
4.1.1. Enumeration of Contraction Paths
Let the number of input tensors in the SpTTN be $n$. To enumerate contraction paths, we employ a function to pick and contract all combinations of two tensors from the list of input tensors. We then recurse over a new list constructed by replacing the contracted tensors with its output. This approach has been studied in the context of finding an optimal contraction sequence for dense tensor networks (Pfeifer et al., 2014). The cost can be analyzed by the recurrence relation, $T(n)={n\choose 2}\cdot T(n-1)$ and $T(2)=1$ (when there are two tensors to contract). The number of valid contraction paths for $n$ tensors is $O\left(\frac{(n!)^{2}}{n\cdot 2^{n}}\right)$.
4.1.2. Enumeration of index orders for a given contraction path
For a given contraction path, we construct a path graph for each term by picking an index order for that term. The path graphs are then fused to obtain a fully fused loop nest tree. Each choice of index order yields a different fused loop nest.
Let the set of indices in the $i$th term be $I_{i}$. The set of indices in the SpTTN is given by $I=\bigcup_{i=1}^{n-1}I_{i}$. We do an exhaustive search by enumerating all index orders independently for each path graph and then considering all possible combinations of these orders. Since we do not allow any repeated indices in our path graphs, the loop nests generated in such an enumeration are unique and span all the possible loop nests for a given contraction path. The cardinality of this exhaustive search is given by $\prod_{i=1}^{n-1}|I_{i}|!$. We later restrict the search to only those index orders that respect the order of the sparse tensor, so if a term involves $k$ sparse indices, the number of possible orders for the term is only $|I_{i}|!/k!$.
4.1.3. Enumeration of Loop Nests
For a given SpTTN, the number of loop nests we enumerate has an upper bound given by the product of the number of contraction paths and the number of index orders for a given contraction path, i.e., $O\Big{(}\frac{(n!)^{2}\cdot\prod_{i=1}^{n-1}|I_{i}|!}{n\cdot 2^{n}}\Big{)}$.
In the following section, we present a dynamic programming algorithm to prune the search space for index order enumeration.
4.2. Finding Cost-Optimal Loop Nests
For a given contraction path, we propose a dynamic programming approach to pick an index order that minimizes a chosen cost metric. Given a fixed contraction path order (or a subsequence of the terms, which defines a subproblem), we seek to find a loop nest tree that minimizes cost. Analogously, in (Ibrahim et al., 2022), dynamic programming is used to find the cost optimal contraction path (tree) given a fixed order of tensors to be contracted [^1] While the two approaches use a similar algorithmic technique, they target different problems and are complementary..
4.2.1. Formal Definition of Contraction Path and Loop Nest
In order to provide a general definition of the algorithm and prove its correctness, we first provide a formal definition of the input (a contraction path) and output (a fully fused loop nest tree) of the algorithm.

### v2

@@ Abstract.
Sparse tensor decomposition and completion are common in numerous applications, ranging from machine learning to computational quantum chemistry. Typically, the main bottleneck in optimization of these models are contractions of a single large sparse tensor with a network of several dense matrices or tensors (SpTTN). Prior works on high-performance tensor decomposition and completion have focused on performance and scalability optimizations for specific SpTTN kernels. We present algorithms and a runtime system for identifying and executing the most efficient loop nest for any SpTTN kernel. We consider both enumeration of such loop nests for autotuning and efficient algorithms for finding the lowest cost loop nest for simpler metrics, such as buffer size or cache miss models. Our runtime system identifies the best choice of loop nest without user guidance, and also provides a distributed-memory parallelization of SpTTN kernels. We evaluate our framework using both real-world and synthetic tensors. Our results demonstrate that our approach outperforms available generalized state-of-the-art libraries and matches the performance of specialized codes.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 36th ACM Symposium on Parallelism in Algorithms and Architectures; June 17–21, 2024; Nantes, Francebooktitle: Proceedings of the 36th ACM Symposium on Parallelism in Algorithms and Architectures (SPAA ’24), June 17–21, 2024, Nantes, Francedoi: 10.1145/3626183.3659985isbn: 979-8-4007-0416-1/24/06ccs: Theory of computation Dynamic programmingccs: Theory of computation Massively parallel algorithmsccs: Software and its engineering Runtime environments
@@ 1. Introduction
Prior works with a focus on high-performance tensor decomposition and completion have introduced efficient and parallel implementations for many SpTTN kernels (Smith et al., 2015; Kaya and Uçar, 2015; Li et al., 2017; Phipps and Kolda, 2019; Choi et al., 2018; Li et al., 2018). Most of these works focus specifically on one or two kernels needed for a particular tensor decomposition, e.g., the matricized Khatri-Rao product (MTTKRP) for CP decomposition (Choi and Vishwanathan, 2014; Kang et al., 2012; Blanco et al., 2018) or the tensor times matrix chain (TTMc) kernel for Tucker (Smith and Karypis, 2017; Oh et al., 2018)). Even for a single decomposition, different algorithms often rely on different SpTTN kernels (Singh et al., 2022). By developing algorithms and libraries for arbitrary SpTTN kernels, we provide functionality for contraction arising (e.g., as a result of a gradient calculation, described in Section 3) from any decomposition/network consisting of dense tensors.
The main challenge in implementation of an SpTTN kernel is finding the most efficient loop nest. In line with prior work (Kjolstad et al., 2017), we represent such loop nest as a tree, in which each vertex is a loop and its descendants are the loops contained within it. In Section 4, we show how to enumerate all loop nests (assuming fusion is done wherever possible) for a given SpTTN. Since each loop order for any pair of contracted tensors yields a distinct loop nest, the size of this space grows factorially in the loop nest depth $m$ and exponentially in the number of tensors $N$. We provide a dynamic programming algorithm to find a cost-optimal loop nest with substantially lower cost, namely $O(N^{3}2^{m}m)$ instead of $O((m!)^{N})$. We state the algorithm for a general cost function that can be decomposed according to the loop nest tree structure, then provide specific cost functions to minimize buffer size and cache misses.
@@ 2. Background
(b) Vertices $i$ and $j$ are fused across the two pairwise contractions.
(c) Vertices $i$, $j$ and $s$ are fused across the two pairwise contractions.
(d) None of the vertices can be fused.
Figure 1. Graphs illustrating loop nests for computing an order 3 TTMc kernel. Sparse loops are shown as dotted vertices.
@@ 9 $\mathcal{S}$[i,r,s] += $\mathcal{X}$ $\mathcal{U}$[j,r]
Listing 3: TTMc kernel computed using the factorize-and-fuse approach. Indices $i$,$j$,$s$ are fused.
2.1. Tensor Notation
We use calligraphic letters to denote tensors, e.g., $\mathcal{T}$. An order $N$ tensor corresponds to an $N$-dimensional array. We denote elements of tensors in parenthesis, e.g., $\mathcal{T}(i,j,k,l)$ for an order 4 tensor $\mathcal{T}$. The indices that do not appear in the output tensor are considered to be summed (contracted). We use capitalized letters to denote the dimensions of the respective indices. For example, the dimension of index $i$ in $\mathcal{A}(i,j)$ is denoted as $I$.
2.2. Tensor Sparsity and Sparse Storage
We use one of the most common ways to store sparse tensors, the Compressed Sparse Fiber (CSF) format (Smith and Karypis, 2015). We refer to the total number of nonzero elements of a tensor $\mathcal{T}$ as $\operatorname{nnz}(\mathcal{T})$. For a sparse tensor $\mathcal{T}$ with $d$ dimensions of size $I_{1},\ldots,I_{d}$, we represent the number of non-zeroes in the $k$th level of the CSF tree for $\mathcal{T}$ (with the first index being at the root) as $\operatorname{nnz}^{(I_{1}\cdots I_{k})}(\mathcal{T})$. Equivalently, this nonzero count may be obtained by considering the number of nonzeros in a reduced tensor obtained by summing away the remaining modes, i.e., $\operatorname{nnz}^{(I_{1}\cdots I_{k})}(\mathcal{T})=\operatorname{nnz}(\mathcal{S})$, where $\mathcal{S}(i_{1},\ldots,i_{k})=\sum_{i_{k+1},\ldots,i_{d}}|\mathcal{T}(i_{1},\ldots,i_{d})|$.
2.3. Tensor Decomposition and Completion Algorithms
Tensor decomposition (Kolda and Bader, 2009) and completion (Smith et al., 2016) refer to the problem of decomposing a tensor into a combination of smaller tensors and estimating missing or incomplete values in a tensor, respectively.
2.4.1. Unfactorized Contraction
To compute a kernel, we can iterate over all the indices and simultaneously contract all the input tensors in the innermost loop. We refer to this approach as unfactorized. This unfactorized loop nest has a depth equal to the number of unique indices. For example, consider an order 3 TTMc kernel in Equation 2. The number of operations is $3\operatorname{nnz}(\mathcal{T})\cdot R\cdot S$ to leading order. In compiler driven frameworks such as TACO (Kjolstad et al., 2017) and COMET (Tian et al., 2021), the schedule generated by default is unfactorized.
The unfactorized approach is optimal in cost for computing certain kernels. For example, the MTTKRP kernel in Equation 1 can be computed using the unfactorized approach with an optimal loop depth of 4. But this approach is asymptotically suboptimal for many other kernels, such as the TTMc.
2.4.2. Pairwise Contraction
A kernel can be computed with minimal asymptotic complexity (loop depth) by contracting the input tensors pairwise. We refer to this approach as pairwise contraction. It is typically used in libraries designed for dense tensor contractions, such as CTF (Solomonik et al., 2014), Tensor Computation Engine (TCE) (Baumgartner et al., 2005), and DEinsum (Ziogas et al., 2022). For example, consider the TTMc kernel in Equation 2. One way in which the tensors can be contracted pairwise is to first contract $\mathcal{T}$ with $\mathcal{V}$, and then its result with $\mathcal{U}$. Each pairwise contraction has an independent loop nest as shown in Listing LABEL:lst:ttmc_pairwise. Both the loop nests have a depth of 4, and the computational cost is $2\operatorname{nnz}(\mathcal{T})\cdot S+2\operatorname{nnz}^{(IJ)}\cdot S\cdot R$ to leading order. Even though an unfactorized approach for computing the MTTKRP kernel (Equation 1) has an optimal loop depth, up to a third of the operations can be eliminated by using pairwise contraction. The unfactorized approach requires $3\cdot\operatorname{nnz}(\mathcal{T})\cdot A$ scalar operations, while the pairwise approach requires $2\operatorname{nnz}^{(IJK)}(\mathcal{T})\cdot A+2\operatorname{nnz}^{(IJ)}\cdot A$ operations.
For contractions involving only dense tensors, the pairwise approach can provide an optimal schedule. But for sparse tensors, whose dimensions are often large, this approach can lead to unmanageable memory requirements for storing dense intermediate tensors. In practice, pairwise contraction with sparse storage of such an intermediates has been observed to be much slower than hand-tuned or even unfactorized implementations for SpTTN kernels (Singh et al., 2022).
2.4.3. Factorize-and-Fuse
[…]
3. SpTTN Kernels
The optimization methods generally leverage all or parts of the gradient of the residual error norm, which yields a contraction of the sparse tensor, with subsets (all but one of the) tensors from the decomposition. The terms involving $\mathcal{T}$ when computing the gradient of $f$ are cost-dominant. Similarly, when computing the residual error ($\rho$) for tensor completion, which is often employed, e.g., in coordinate descent methods, the terms involving the sparse tensor are cost-dominant. $\rho=\mathcal{T}-\Omega\ast\tilde{\mathcal{T}}(A_{1},\ldots,A_{n})$, where $\ast$ is the Hadamard (pointwise product), the sparse tensor $\Omega$ represents the set of observed entries in the input tensor $\mathcal{T}$ and $\tilde{\mathcal{S}}$ is the output tensor obtained by contracting $\Omega$ with a network of dense factors.
[…]
3.1. Loop Nests and Loop Nest Forests
The computation of a tensor contraction generally involves loop nests of some form. Any loop nest can be represented by an ordered tree or forest, each vertex of which is a loop, and its ordered children are the loop nests contained directly in that loop. Each leaf corresponds to a contraction term (a pair of tensors contracted together). For example, the loop nest in Listing LABEL:lst:ttmc_pairwise is represented by the tree in Figure 1(a). We refer to a tree with a single leaf as a path graph. A similar representation is used in TACO (Kjolstad et al., 2017).
The leaves of the loop nest tree define the order in which the contraction terms are executed. We refer to this order as the contraction path. A contraction path for a kernel is valid if we can obtain the output tensor by executing the contraction terms in the order specified by it.